# Multi-messenger lensing time delay as a probe of the graviton mass

### v1

@@ IV.3 Total time delay
$\Delta t_{\rm g}$ $=\left(1+\frac{m^{2}}{2\omega^{2}}\right)(1+z_{l})\frac{D_{s}D_{l}}{D_{ls}}\hat{\phi}(\boldsymbol{\theta}_{\rm g}^{\rm I},\boldsymbol{\beta})$ (38)
@@
Note that Eq. (38) confirms the intuition of [39]. To make connection with the electromagnetic time-delay, one must acknowledge that the images form at different angles $\boldsymbol{\theta}_{\rm g}^{\rm I}$ than in the massless case $\boldsymbol{\theta}_{\gamma}^{\rm I}$, as we have found in Sec. III. This comes from the fact that massive gravitons are more deflected than massless photons such that the geometric time delay is longer. However, since they hit the lens plane with a larger impact parameter, they also experience less Shapiro time-delay. Expanding $\boldsymbol{\theta}_{\rm g}^{\rm I}$ around $\boldsymbol{\theta}_{\gamma}^{\rm I}$ to first order in $m^{2}/\omega^{2}$, one obtains that the massive GW time delay $\Delta t_{\rm g}$ relates to their massless counterpart $\Delta t_{\gamma}$ as
@@
where the electromagnetic time delay $\Delta t_{\gamma}$ can be defined in terms of the standard time delay (Eq. 41)
@@
The experienced lensing reader might recognize the lensing equation in the square brackets of Eq. (43), which tell that EM images form at positions $\boldsymbol{\theta}_{\gamma}^{\rm I}$, which extremise the Fermat potential
@@
Eq. (46) is the main result of this article. It relates the time delay between different images in the GW and EM sector if GWs obey a massive dispersion relation, as in Eq. (5). The interpretation is clear, massive gravitons travel slower than massless photons over the path difference of different images. It is quite remarkable that even if they travel through different paths than massless particles, the amount of time gained geometrically is canceled by experiencing less Shapiro time delay. We depict this in Fig. 2. Also note that while we computed the scattering angle for a point-like lens in Sec. III, this cancellation is completely general and holds for arbitrary lens models. This fact has been underappreciated in the literature [34], where it was assumed that in modified gravity, GWs traveling at $c_{\rm g}<1$ also travel along the same paths, although this is generically incorrect as shown in Sec. III. It had to be similarly assumed that the Shapiro time delay is unaffected by modified gravity. We showed here instead that if the GWs follow different paths than photons, the two contributions individually differ from the massless case and they cancel in the time delay.
@@
We now discuss the observational prospects. Suppose that we observe a multi-messenger event. In a golden scenario, one can imagine observing a strongly lensed gravitational wave in coincidence with an electromagnetic counterpart. This would lead to a time delay measurement in the electromagnetic domain ($\Delta t_{\gamma}$) and one in the gravitational wave ($\Delta t_{\rm g}$) domain. These should be related as in Eq.(46). Rearranging this formula, the mass of the graviton can be expressed as
@@
where we have temporarily introduced factors of $\hbar=1=c$. The advantage of the multi-messenger measurement is quite clear. It does not require any modeling of the lens or of cosmology, which disappears in the comparison between the electromagnetic and GW time delay. This is quite remarkable, given that they follow different geodesics and experience different local Shapiro time delays. We may even get a different time delay for each frequency $\Delta t_{\rm g}(\omega)$ if we can identify corresponding crests of the GW for the different signals. Note that the advantage with respect to dispersion relation constraints on the speed of the low versus high frequencies of the GW, is that it does not require us to estimate the GW phase at the source, which requires a waveform model in the presence of a massive graviton. Here instead, one can identify the merger time and the peaks of the GW in a model independent way. This allows for a fully model-independent constraint on the graviton mass. Hereafter, we evaluate whether this constraint is competitive.
In the null mass case, the time delays are equal, such that $t\equiv\Delta t_{\rm g}-\Delta t_{\gamma}=0$ up to some uncertainty $\sigma_{t}=\sqrt{\sigma_{\Delta t_{\rm g}}^{2}+\sigma_{\Delta t_{\gamma}}^{2}}$. The $95\%$ confidence limit on $m$ reads
@@
The uncertainty on the merger time of the gravitational wave is determined by how well one can resolve the waveform. We adopt a time delay error in the GW signal of $0.1$s, which is predicted to be feasible for high signal to noise ratio events with LISA [49, 50]. For the uncertainty on the electromagnetic time delay, we consider two scenarios. In an optimistic case, the time delay can be determined to sub-second precision $\sigma_{\Delta t_{\gamma}}=0.1$s, as might be the case for a short gamma ray burst [51]. A more pessimistic EM time delay uncertainty could be $\sigma_{\Delta t_{\gamma}}=10^{5}$s (similar to the typical kilonova time delay uncertainty [52]). We plot the $2\sigma$ upper limit on $m$ as a function of frequency $f$ (where $f=\omega/(2\pi)$) for these two scenarios for a signal with $\Delta t_{\gamma}=1000$ days in Fig. 3. Note that this time delay is typical for galactic cluster lenses and that galaxy lenses would typically give lower time delays, and hence weaker constraints on the graviton mass. The tightest constraint shown is $m<3\cdot 10^{-23}$eV/c2 for one measurement of $\Delta t_{\gamma}$ and $\Delta t_{\rm g}$ (i.e. two images detected) in the LISA band, with $f=10^{-4}$. If we were to detect a quadruply-imaged signal, we would obtain three independent measurements of both time delays which implies $m<2\cdot 10^{-23}$eV/c2. This constraint is comparable to current dispersion relation bounds [40] with the adventage of being independent of the lens, waveform, and cosmological models. Additionally, one could in principle measure the time delay at different frequencies if one can identify precisely waveform features of the different images, which would provide additional independent measurements of $\Delta t_{\rm g}$.
@@
Figure 3: Upper bound on the graviton mass as a function of gravitational wave frequency. The solid line corresponds to an optimistic scenario with error on EM time delay measurements of $0.1$s, while the dashed line corresponds to an error of $10^{5}$s. The gray areas above the lines correspond to the values of $m$ excluded at $95\%$ confidence. The area shaded in blue corresponds to the LISA frequency band, while the green area corresponds to ground-based detector bands.
@@ V Magnification constraint
In this section, we derive the amplification factor of gravitational waves in presence of a non-zero mass in the dispersion relation (Eq. 5). We first show that the Kirchhoff diffraction integral is unaffected by the graviton mass in Sec. V.1. We then compute the amplification factor in Sec. V.2. Finally, we show in Sec. V that the comparison of EM versus GW magnification leads to worse constraints on the graviton mass than the time delay technique.
@@ V.1 Kirchhoff’s theorem
Figure 4: Illustration of the Kirchhoff theorem setup. We consider a volume $V$ in blue bounded by a surface $S$ in order to evaluate the wave amplitude at point $\boldsymbol{x}_{o}$, which corresponds to the observer. The source is located outside the volume $V$. We also depict a sphere of small radius $\epsilon$ around the observer (the excluded region in our calculation) and the vector $\boldsymbol{n}$, which is normal to the surface $S$. Additionally, though the Kirchhoff theorem is general, we show the positions of a source, observer and lens together with the paths followed by the geometric optics images which form at impact parameters $\boldsymbol{b}_{\rm G}(\boldsymbol{\theta}^{\rm{I}}_{\rm g})$ and $\boldsymbol{b}_{\rm G}(\boldsymbol{\theta}^{\rm{J}}_{\rm g})$ in the lens plane to facilitate the connection between our derivation of the theorem and its use in the context of lensing.
To evaluate how the amplification factor of a lensed gravitational wave differs in the presence of a graviton mass, we start by demonstrating that Kirchhoff’s theorem holds unchanged. This is true despite the modified dispersion relation. This theorem allows one to express the amplitude of the wave at a point in terms of the wave evaluated on a closed surface. This is useful when one knows the amplitude of a wave on that surface but not at the point in question, because of intervening matter between the source and the observer, as is the case in a gravitational lensing situation. Following the derivation in chapter 8.3 of Ref. [53], we assume that the GW is well described by a scalar wave, which we write in Fourier space [^5] Note that we can treat GWs as scalar waves, since to lowest order in geometric optics, the polarization is parallel transported along the geodesic [54], but see [44, 45] for beyond geometric optics effects.
@@
on Minkowski space. Suppose that $\tilde{U}^{\prime}$ is another solution of Eq. (50). If $\boldsymbol{n}$ is the inward normal to a closed surface $S$, and $\tilde{U}$ and $\tilde{U}^{\prime}$’s first and second partial derivatives are continuous within and on $S$, we can apply Green’s theorem over the enclosed volume $V$
@@
Substituting Eq. (50) into the left hand side, the integral vanishes in the same way it does in the massless case, because scalars commute. We are left with
@@ V.2 Amplification factor
Eq. (60) suggests that one can use the comparison of the GW and EM amplifications to constrain the graviton mass. [^6] The EM amplification is not directly accessible because one observes fluxes instead of the amplitude of the EM wave, which means one has to resort to magnifications, as discussed in the next section. Note that this assumes that the source angles are the same, i.e. $\boldsymbol{\beta}_{\rm g}=\boldsymbol{\beta}_{\gamma}$. This might be the case for an EM transient emitted along the GW signal. For an extended source, such as a galaxy, it becomes unclear where the GW signal originates from within the host galaxy, such that in general, $\boldsymbol{\beta}_{\rm g}\neq\boldsymbol{\beta}_{\gamma}$. The difference in amplification factor between a GW being emitted from one side of the galaxy and the other might be larger than the difference generated by the mass of the graviton. This would of course spoil the test. In the following, we show that even in the best scenario, this method to constrain the graviton mass is weaker and less practical than the time delay technique, as it requires assumptions about the lens model and cosmology
@@ V.3 Magnification
$\mu_{\rm g}=\left(1-\frac{m^{2}}{\omega^{2}}\right)\mu_{\gamma}\,$, (65)
Eq. (65) suggests that one can use magnifications to constrain the graviton mass. We consider again the golden scenario, where one observes a lensed GW and the lensed host galaxy, both of which are magnified according to Eq.(63) and (64) with $\mu^{\rm{I}}_{g,\gamma}$ with $\rm{I}\in\{1,\dots,N\}$ for $N$ images. While we have $2N$ measurements ($N$ $h^{\rm{I}}$ for the GW images and NN $\Phi^{\rm{I}}$ for the EM images), we need extra information to break the degeneracy between $h_{\rm nolens}$, $\Phi_{\rm nolens}$ and the magnifications $\mu_{\rm g}^{\rm{I}}$ and $\mu_{\gamma}^{\rm{I}}$ which are unobservable and represent $2N+2$ unknowns. The EM magnification can be estimated from a lens model [56], for which current error bars are of the order of $20\%$ due to the mass-sheet degeneracy [57, 58]. The GW magnification can be extracted from the luminosity distance encoded in the waveform, the source redshift and a cosmological model, which allow to compute $\bar{D}_{\rm L}(z)$. In this case the current error bars can be of the order of $30\%$, driven by errors in the observed luminosity distance which are largely due to its degeneracy with inclination. Under these assumptions, the inferred magnifications of each image can be used to extract the graviton mass which reads
@@
$m$ $<\frac{\sqrt{2}\hbar\omega}{c^{2}}\sqrt[4]{\left(\frac{\sigma_{\mu_{\rm g}}}{\mu_{\rm g}}\right)^{2}+\left(\frac{\sigma_{\mu_{\gamma}}}{\mu_{\gamma}}\right)^{2}}$ (67)
@@
Even under the most optimistic scenario – assuming a perfect lens model, known cosmology, and minimal GW luminosity distance error (e.g. from a particularly loud event or broken distance-inclination degeneracy) – yielding relative error bars on the magnifications of around $1\%$, this constraint remains at least 3 orders of magnitude weaker than the one from the time delay. It also requires modeling the lens to extract the EM magnifications $\mu_{\gamma}^{\rm{I}}$ and to assume a cosmological model to extract the GW magnification $\mu^{\rm{I}}_{g}$. Furthermore, we had to assume that the source angle in the EM and GW are the same, which may not be a good assumption if the EM source is extended, like a galaxy. In all regards, this constraint is weaker than the constraint coming from the time delay.

### v2

@@ IV.3 Total time delay
$\Delta t_{\rm g}$ $=\left(1+\frac{m^{2}}{2\omega^{2}}\right)(1+z_{l})\frac{D_{s}D_{l}}{D_{ls}}\hat{\phi}(\boldsymbol{\theta}_{\rm g}^{\rm I},\boldsymbol{\beta})$
@@
Note that Eq. (35) confirms the intuition of [21]. To make connection with the electromagnetic time delay, one must acknowledge that the images form at different angles $\boldsymbol{\theta}_{\rm g}^{\rm I}$ than in the massless case $\boldsymbol{\theta}_{\gamma}^{\rm I}$, as we have found in Sec. III. This comes from the fact that massive gravitons are more deflected than massless photons such that the geometric time delay is longer. However, since they hit the lens plane with a larger impact parameter, they also experience less Shapiro time delay. Expanding $\boldsymbol{\theta}_{\rm g}^{\rm I}$ around $\boldsymbol{\theta}_{\gamma}^{\rm I}$ to first order in $m^{2}/\omega^{2}$, one obtains that the massive GW time delay $\Delta t_{\rm g}$ relates to their massless counterpart $\Delta t_{\gamma}$ as
@@
where the electromagnetic time delay $\Delta t_{\gamma}$ can be defined in terms of the standard time delay [Eq. (37)]
@@
The experienced lensing reader might recognize the lensing equation in the square brackets of Eq. (39), which tell that EM images form at positions $\boldsymbol{\theta}_{\gamma}^{\rm I}$, which extremize the Fermat potential
@@
Equation (42) is the main result of this article. It relates the time delay between different images in the GW and EM sector if GWs obey a massive dispersion relation, as in Eq. (6). The interpretation is clear, massive gravitons travel slower than massless photons over the path difference of different images. It is quite remarkable that even if they travel through different paths than massless particles, the amount of time gained geometrically is canceled by experiencing less Shapiro time delay. We depict this in Fig. 2. Also note that while we computed the scattering angle for a pointlike lens in Sec. III, this cancellation is completely general and holds for arbitrary lens models. This fact has been underappreciated in the literature [23], where it was assumed that in modified gravity, GWs traveling at $c_{\rm g}<1$ also travel along the same paths, although this is generically incorrect as shown in Sec. III. It had to be similarly assumed that the Shapiro time delay is unaffected by modified gravity. We showed here instead that if the GWs follow different paths than photons, the two contributions individually differ from the massless case and they cancel in the time delay.
@@
We now discuss the observational prospects. Suppose that we observe a multimessenger event. In a golden scenario, one can imagine observing a strongly lensed gravitational wave in coincidence with an electromagnetic counterpart. This would lead to a time-delay measurement in the electromagnetic domain ($\Delta t_{\gamma}$) and one in the gravitational wave ($\Delta t_{\rm g}$) domain. These should be related as in Eq.(42). Rearranging this formula, the mass of the graviton can be expressed as
@@
where we have temporarily introduced factors of $\hbar=1=c$. The advantage of the multimessenger measurement is quite clear. It does not require any modeling of the lens or of cosmology, which disappears in the comparison between the electromagnetic and GW time delay. This is quite remarkable, given that they follow different geodesics and experience different local Shapiro time delays. We may even get a different time delay for each frequency $\Delta t_{\rm g}(\omega)$ if we can identify corresponding crests of the GW for the different signals.
Note that the advantage with respect to dispersion relation constraints on the speed of the low versus high frequencies of the GW, is that it does not require us to estimate the GW phase at the source, which requires a waveform model in the presence of a massive graviton. Here instead, one can identify the merger time and the peaks of the GW in a model independent way. This allows for a fully model-independent constraint on the graviton mass. Hereafter, we evaluate whether this constraint is competitive.
In the null mass case, the time delays are equal, such that $t\equiv\Delta t_{\rm g}-\Delta t_{\gamma}=0$ up to some uncertainty $\sigma_{\Delta t}=\sqrt{\sigma_{\Delta t_{\rm g}}^{2}+\sigma_{\Delta t_{\gamma}}^{2}}$. For one time-delay measurement, we take two time-of-arrival measurements, whose errors $\sigma_{t}$ are added in quadrature.  [^5] Hence $\sigma_{\Delta t}=\sqrt{2}\sigma_{t}$ for both GW and electromagnetic signals. The $95\%$ confidence limit on $m$ reads
@@
Note that the individual contributions of geometric and Shapiro effects to the total time delay in the denominator are degenerate. The degeneracy is broken by the details of the lens model, which drops out of this constraint. As such, the constraint holds regardless of the precise cancellation between the Shapiro and geometric time delays.
The uncertainty on the merger time of the gravitational wave is determined by how well one can resolve the waveform.
We adopt a time-delay error in the GW signal of $10^{-6}$ for the LIGO band, and $0.1$s for LISA, which is predicted to be feasible for high signal to noise ratio events [52, 62]. This distinction is not apparent in the plot as the term in the square root is dominated by the EM time-delay errors.
@@
For the uncertainty on the electromagnetic time delay, we consider two scenarios. In an optimistic case, the time delay can be determined to subsecond precision $\sigma_{\Delta t_{\gamma}}=0.1$s, as might be the case for a short gamma ray burst [1]. A more pessimistic EM time-delay uncertainty could be $\sigma_{\Delta t_{\gamma}}=10^{5}$ s (similar to the typical kilonova time-delay uncertainty [19]). We plot the $2\sigma$ upper limit on $m$ as a function of frequency $f$ (where $f=\omega/(2\pi)$) for these two scenarios for a signal with $\Delta t_{\gamma}=1000$ days in Fig. 3. Note that this time delay is typical for galactic cluster lenses and that galaxy lenses would typically give lower time delays and, hence, weaker constraints on the graviton mass. The tightest constraint shown is $m<3\cdot 10^{-23}$eV/c2 for one measurement of $\Delta t_{\gamma}$ and $\Delta t_{\rm g}$ (i.e. two images detected) in the LISA band, with $f=10^{-4}$Hz lensed by a cluster. If we were to detect a quadruply imaged signal, we would obtain three independent measurements of both time delays which implies $m<2\cdot 10^{-23}$eV/c2. This constraint is comparable to current dispersion relation bounds [2] with the advantage of being independent of the lens, waveform, and cosmological models. Additionally, one could in principle measure the time delay at different frequencies if one can identify precisely waveform features of the different images, which would provide additional independent measurements of $\Delta t_{\rm g}$.
As mentioned above, we are working in the geometric optics approximation, which may break down when the wavelength of the GW becomes smaller or comparable to the Schwarzschild radius of the lens, $\lambda\approx R_{\rm S}$. This may happen at low frequencies for certain galaxy-scale lenses, in which case the formalism that we use is unsuitable, and one should resort to wave optics. [^6] An example case where geometric optics breaks down occurs at frequencies around $10^{-7}$Hz, which is much lower than LISA merger frequency. At this frequency the GW wavelength is $\lambda\approx 10^{15}$m, comparable to a Schwarzschild radius of a galaxy similar to the Milky Way. For galactic cluster lenses, with $R_{\rm S}$ of the order of light years, this is at least 4 orders of magnitude larger than a GW wavelength at the lowest frequency we consider ($10^{-4}$Hz, which corresponds to roughly $10^{12}$ m). In this scenario, the quoted constraints in Fig. 3 are safely within the geometric optics regime.
@@
Figure 3: Upper bound on the graviton mass as a function of gravitational wave frequency. The solid line corresponds to an optimistic scenario with an error on EM time-delay measurements of $0.1$s, while the dashed line corresponds to an error of $10^{5}$s. The gray areas above the lines correspond to the values of $m$ excluded at $95\%$ confidence. The area shaded in blue corresponds to the LISA frequency band, while the green area corresponds to ground-based detector bands. Constraints shown are for a galactic cluster lens.
@@ V Magnification constraint
In this section, we derive the amplification factor of gravitational waves in the presence of a nonzero mass in the dispersion relation [Eq. (6)]. We first show that the Kirchhoff diffraction integral is unaffected by the graviton mass in Sec. V.1. We then compute the amplification factor in Sec. V.2. Finally, we show in Sec. V.3 that the comparison of EM versus GW magnification leads to worse constraints on the graviton mass than the time delay technique.
@@ V.1 Kirchhoff’s theorem
Figure 4: The Kirchhoff theorem setup. We consider a volume $V$ in blue bounded by a surface $S$ to evaluate the wave amplitude at point $\boldsymbol{x}_{o}$, which corresponds to the observer. The source is located outside the volume $V$. We also depict a sphere of small radius $\epsilon$ around the observer (the excluded region in our calculation) and the vector $\boldsymbol{n}$, which is normal to the surface $S$. Additionally, though the Kirchhoff theorem is general, we show the positions of a source, observer and lens together with the paths followed by the geometric optics images which form at impact parameters $\boldsymbol{b}_{\rm G}(\boldsymbol{\theta}^{\rm{I}}_{\rm g})$ and $\boldsymbol{b}_{\rm G}(\boldsymbol{\theta}^{\rm{J}}_{\rm g})$ in the lens plane to facilitate the connection between our derivation of the theorem and its use in the context of lensing.
To evaluate how the amplification factor of a lensed gravitational wave differs in the presence of a graviton mass, we start by demonstrating that Kirchhoff’s theorem holds unchanged. This is true despite the modified dispersion relation. This theorem allows one to express the amplitude of the wave at a point in terms of the wave evaluated on a closed surface. This is useful when one knows the amplitude of a wave on that surface but not at the point in question, because of intervening matter between the source and the observer, as is the case in a gravitational lensing situation. Following the derivation in Chap. 8.3 of Ref. [18], we assume that the GW is well described by a scalar wave, which we write in Fourier space [^7] Note that we can treat GWs as scalar waves, since to lowest order in geometric optics, the polarization is parallel transported along the geodesic [55], but see [26, 27] for beyond geometric optics effects.
@@
on Minkowski space. Suppose that $\tilde{U}^{\prime}$ is another solution of Eq. (46). Following the setup showin in Fig.  4, if $\boldsymbol{n}$ is the inward normal to a closed surface $S$, and $\tilde{U}$ and $\tilde{U}^{\prime}$’s first and second partial derivatives are continuous within and on $S$, we can apply Green’s theorem over the enclosed volume $V$
@@
Substituting Eq. (46) into the left-hand side, the integral vanishes in the same way it does in the massless case, because scalars commute. We are left with
@@ V.2 Amplification factor
Equation (56) suggests that one can use the comparison of the GW and EM amplifications to constrain the graviton mass. [^8] The EM amplification is not directly accessible because one observes fluxes instead of the amplitude of the EM wave, which means one has to resort to magnifications, as discussed in the next section. Note that this assumes that the source angles are the same, i.e. $\boldsymbol{\beta}_{\rm g}=\boldsymbol{\beta}_{\gamma}$. This might be the case for an EM transient emitted along the GW signal. For an extended source, such as a galaxy, it becomes unclear where the GW signal originates from within the host galaxy, such that in general, $\boldsymbol{\beta}_{\rm g}\neq\boldsymbol{\beta}_{\gamma}$. The difference in amplification factor between a GW being emitted from one side of the galaxy and the other might be larger than the difference generated by the mass of the graviton. This would of course spoil the test. In the following, we show that even in the best scenario, this method to constrain the graviton mass is weaker and less practical than the time-delay technique, as it requires assumptions about the lens model and cosmology
@@ V.3 Magnification
$\mu_{\rm g}=\left(1-\frac{m^{2}}{\omega^{2}}\right)\mu_{\gamma}\,.$ (61)
Equation (61) suggests that one can use magnifications to constrain the graviton mass. We consider again the golden scenario, where one observes a lensed GW and the lensed host galaxy, both of which are magnified according to Eqs.(59) and (60) with $\mu^{\rm{I}}_{g,\gamma}$ with $\rm{I}\in\{1,\dots,N\}$ for $N$ images. While we have $2N$ measurements ($N$ $h^{\rm{I}}$ for the GW images and NN $\Phi^{\rm{I}}$ for the EM images), we need extra information to break the degeneracy between $h_{\rm nolens}$, $\Phi_{\rm nolens}$ and the magnifications $\mu_{\rm g}^{\rm{I}}$ and $\mu_{\gamma}^{\rm{I}}$ which are unobservable and represent $2N+2$ unknowns. The EM magnification can be estimated from a lens model [60], for which current error bars are of the order of $20\%$ due to the mass-sheet degeneracy [15, 42]. The GW magnification can be extracted from the luminosity distance encoded in the waveform, the source redshift and a cosmological model, which allow one to compute $\bar{D}_{\rm L}(z)$. In this case the current error bars can be of the order of $30\%$, driven by errors in the observed luminosity distance which are largely due to its degeneracy with inclination. Under these assumptions, the inferred magnifications of each image can be used to extract the graviton mass which reads
@@
$m$ $<\frac{\sqrt{2}\hbar\omega}{c^{2}}\sqrt[4]{\left(\frac{\sigma_{\mu_{\rm g}}}{\mu_{\rm g}}\right)^{2}+\left(\frac{\sigma_{\mu_{\gamma}}}{\mu_{\gamma}}\right)^{2}}$
@@
Even under the most optimistic scenario – assuming a perfect lens model, known cosmology, and minimal GW luminosity distance error (e.g. from a particularly loud event or broken distance-inclination degeneracy) – yielding relative error bars on the magnifications of around $1\%$, this constraint remains at least 3 orders of magnitude weaker than the one from the time delay. It also requires modeling the lens to extract the EM magnifications $\mu_{\gamma}^{\rm{I}}$ and to assume a cosmological model to extract the GW magnification $\mu^{\rm{I}}_{g}$. Furthermore, we had to assume that the source angle in the EM and GW are the same, which may not be a good assumption if the EM source is extended, like a galaxy. Note that if the theory allows for extra polarizations, some power may be lost from the spin-2 channels. This could lead to biases in the GW distance and magnification if this effect is not properly taken into account, which represents further complications. However, testing for extra polarizations can be done in the absence of lensing [2]. In all regards, this constraint is weaker than the constraint coming from the time delay.